OpenStudy (anonymous):

The area of a rectangular pool is 48 square meters. The length is three times as long as the width. What are the dimensions of the pool? A. 2 and 24 B. 3 and 16 C. 4 and 12 D. 6 and 8

OpenStudy (johnweldon1993):

We know that area = length * width \[A = L \times W\] We also know that the length is 3 times the width so \[L = 3W\] So we can substitute that into the first equation for 'L' along with 48 for the area \[48 = 3W \times W\] \[\large 3W^2 = 48\] Can you solve this?

OpenStudy (anonymous):

no :/

OpenStudy (johnweldon1993):

Its okay don't get down on yourself for it \[\large 3W^2 = 48\] We first want to divide both sides by 3 to isolate W² \[\large W^2 = 16\] Now take the square root of both sides to solve for W \[W = \sqrt{16}\] W = ..?

OpenStudy (anonymous):

i get it now (:

OpenStudy (johnweldon1993):

Awesome!

OpenStudy (johnweldon1993):

So thats the width and whats the length?

OpenStudy (anonymous):

8

OpenStudy (johnweldon1993):

Not quite...if we find that W = √16 W = 4 And the length is 3 times that number we have W = 4 L = 12